1.      Perform the indicated operations and simplify if necessary. Convert your answers into (I) decimals and (II) percentages.

a. . 3 − 11+ 9

2    4    8

b.   5 × 21 + 8 ÷ 44

7   55    17    51

 

2.      Perform the indicated operations and convert the results into fractions.

a.   845.23 +1.37− 844.04

b.   12.45%×2.895

 

3.      A rice supplier transports 475 kilograms each of Maharlika, Sinandomeng, Dinorado, and Malagkit to a supermarket every day. The prices per kilogram of the rice variants are P37.50, P38.75, P41.25, and P55.75, respectively.

a.      How much does the supermarket pay the supplier every day?

b.      On a particular day, one of the transport vehicles got into a minor accident and lost some kilograms of Dinorado. The remaining Dinorado was weighed upon delivery, and the supplier was paid in accordance with how much was left. If the total amount paid was P78,375.00, how many kilograms of Dinorado were lost?

4.      A company sells two types of cell phones, C1 and C2. The total number of cell phones sold increased by 60% from 2014 to 2015. Eighty percent of the phones sold in 2015 were of type C2; with the number of units sold (C1 and C2 combined) totaling 152,000.

a.      How many phones were sold in 2015? How many were of type C1?

b.      How many phones were sold in 2014?

c.      Fifty percent of the phones sold in 2014 were of type C1, and the prices of the phones did not change from 2014 to 2015. If each C1 unit costs P10,000.00 and each C2 unit costs P5,000.00, did the company’s revenues increase or decrease? By what percent?


1
Expert's answer
2021-01-24T18:26:00-0500

1a)

3–11+9=1

i) 1.00

ii) 1%

1b)

5×21+8÷44= 1157/11

i) 105.18

iI) 1,(259/5000)%


2a)

845.23+1.37–844.04

=2.56

=64/25

2b)

12.45% × 2.895

= 0.3604

=144172/400000


3a)

The supermarket pays:

(P37.50×475)+(P38.75×475)+(P41.25×475)+(P55.75×475) = P82293.75

3b)

Price of Dinorado: P41.25

Amount paid after the accident: P78375.00

Initial amount paid: P82293.75

P82239.75 – P78375.00=P3918.75

∴ amount of Dinorado lost =3918.75/475

=8.25 kilograms


4a)

Total units sold in 2015 =152000units

a 60% increase in 2015 = 60% × 152000

= 91200 units sold in 2015

ii) 80% of units sold were C2

100% – 80%= 20% (C1)

20% × 91200= 18240 units of C1

4b)

Phone sold in 2014: 152000–91200

=60800 units

50%=C1 ;price of C1=P10000; price of C2=P5000

Profit for C1= sales × price

= P10000 × (50%×60800)

= P304000000

Profit for C2= P5000 × (50%×60800)

= P152000000

Total profit for 2014

=P304000000+P152000000

= P456000000

Profit for 2015

C1=20% ; C2=80% ;units sold=91200

Profit for C2= P5000×(80%×91200)

=P364800000

Proit for C1=P10000×(20%×91200)

=P182400000

Total profit for 2015

=P364800000+P182400000

=P547200000


Difference in revenue 2015 and 2014=P91200000


Revenue decreased by (91200000/456000000)%

=20%